Q) Suppose there are two groups of people in an economy, with their respective consumption function.

                        CA=100+0.5YA

                        CB=150+0.75YB

   planned investment 200.Where Ci and Yi  are respectively consumption and income level

 for the ith group (for all i=A,B). Suppose each group gets 50% of total GDP in the economy.

Determine the equilibrium level of income. Considering the equilibrium income as the initial income, a lumpsum tax T is imposed on Group A and the same amount is given as a transfer to Group B.

 Now determine planned consumption, planned savings , planned aggregate Demand and the Actual Investment as the initial level of income



1
Expert's answer
2021-06-13T15:48:52-0400

(a) Determination of equilibrium level of income

Yi = AE = Ci + I

Ci = CA + CB = 100+ 0.5YA + 150 + 0.75YB

But since each receive 50% of GDP, YA = YB

Yi = AE = 100 + 0.5Y + 150 + 0.75Y + 200 = 450 + 1.25Y


(b) Determination of;

i) Planned consumption

With introduction of lump-sum tax

CA = 100 + 0.5 (YAd) = 100 + 0.5 (Y-T)

CB = 150 + 0.75 (YB+T)


ii) Planned Savings

For group A, Planned comsumption = 100 + 0.5(YA d)

MPC - 0.5, but MPS + MPC = 1

"\therefore" MPS = 1 - 0.5 = 0.5

So planned savings for group B will be 0.5(YA-T).


For group B, planned consumption = 150 + 0.75(YB + T)

MPC = 0.75, but MPS + MPS = 1

"\therefore" MPS = 1 - 0.75 = 0.25

So planned savings for group B = 0.25(YB + T)\


Aggregate Planned Savings = Planned savings for A + Planned savings for B

= 0.5(YA -T) + 0.25(YB+T) = 0.5YA - 0.5T + 0.25YB + 0.25T = 0.5YA + 0.25YB - 0.5T+0.25T

= 0.5YA + 0.25YB -0.25T


iii) Planned AD = Planned Aggregate consumption + Planned Investment

= Planned CA + Planned CB + Planned Investment

= 100 + 0.5 (YA -T) + 150 + 0.75(YB + T) + 200

= 450 + 0.5YA - 0.5T + 0.75YB + 0.75T

= 450 + 0.5YA +0.75YB + 0.25T


iv) At equilibrium, Actual investment (I) = Planned Investment (Ip) = 200